# Visualizationary: Automating Design Feedback for Visualization Designers using LLMs

### v1

@@ 4.3 Visualization Metrics
    Virtual Eyetracker. We utilize the Scanner Deeply [60], a virtual eyetracker that predicts gaze on an overview task of a visualization image. We analyze the results of Scanner Deeply from three perspectives: focus of salience on text [45], center bias [9], and focus on visual attention, represented by low-level stimuli [32, 28].
    • Focus on text. Goal: Detect visualizations that overly focus on text. Metric used: Use OCR to box each letter and compute the rate of salience focused in all of those boxes. We measure the amount of salience focused on text. We use OCR to detect alphabetical letters. Decision: Mark as “overly salient to text” if the salience in text boxes is above the $10^{th}$ percentile of the chart saliency in the dataset by Shin et al. [60].
    • Focus on center. Goal: Detect charts that overly focus on the center (center bias). Metric used: Measure salience in the center. The center is defined as a rectangle in the middle of the image, with its height and width as 1/3×1/3$\times$1 / 3 × of the chart. Decision: Mark as “overly salient in the center” if the center salience is above $10^{th}$ percentile.
    • Focus on visual attention. Goal: Detect visualizations that scarcely focus on visual attention. Metric used: We measure the rate of salience-covered areas of zones near (the radius is 5 pixels) where there is a difference in RGB values, which we call zones with low-level stimuli. Decision: Mark as “scarcely salient” if the salience in zones with low-level stimuli is less than $90^{th}$ percentile.
@@
    Textual characters. We evaluate the existence of textual components in the visualization image [62]. Using text along with charts can help understand visualizations.
    • Title: Goal: If a title is detected, then provide title recommendations. Metric used: Title is detected via DePlot, a VLM that translates visualizations into tables and textual components. Decision: DePlot decides if there is a title in the image.
    • Textual content: Goal: Detect text in visualization image. Metric used: Texts in the image are detected using PyTesseract-OCR. Decision: If the OCR tool detects at least one letter in the text, then it contains the text.
@@
    Visual representations. We observe how data is visually represented. Our focus is on providing chart recommendations based on data table and design preamble, as well as detecting chartjunk, or unnecessary visual embellishments [7, 6] in the design.
    • Optimal chart type: Goal: Given the data, suggest an optimal chart type. Metric used: The data table extracted from DePlot. Decision: ChatGPT makes suggestions based on a preamble of design knowledge (perceptual rankings of visual cues, described in Mackinlay’s APT [41]), and the data table provided by DePlot. To help the designer make an optimal decision, we provide the table as well as the design knowledge instilled in ChatGPT.
    • Visual embellishment (“chart junk”): Goal: Detect chartjunk in the visualization image. Metric used: A computer vision model that detects objects, such as YoloR [67]. Decision: The object detection model detects objects within the chart image.
@@
    Color perception. Detect colors and warn of misusing colors when similar colors are used [23, 50].
    • Variability of colors: Goal: Identify all colors used in the visualization image. Metric used: A customized function created in Python. Decision: The function identifies multiple colors.
    • Similarity of colors: Goal: Identify similar colors in the visualization image. Metric used: A customized function created in Python. A group of similar colors are determined using differences in RGB values between detected colors. Decision: The function identifies a group of similar colors.
@@
    Accessibility. It is also important to create visualizations that are accessible for everyone [1, 20].
    • Color vision deficiency: Goal: Detect visualization images that are hugely affected when simulating the image into how people with color vision deficiency (deuteranopia, protanopia, tritanopia) would see. Metric used: We measure the information loss caused in simulated visualization images, using the difference in entropy loss from its original visualization image. Decision: If the entropy loss is larger than a threshold in simulated visualization images, then, warn of the loss in information for people with color vision deficiency.
We instill these metrics along with the suggestions relevant to the decisions into ChatGPT as a prompt (see Sec. 4.6). Then, based on a series of template questions, ChatGPT provides the appropriate answer.
4.4 Visualization Design Reports
Results from the feedback backend are presented as interactive visualization design reports. A design report is an interactive document consisting of a hierarchy of expandable sections designed in a bottom-up fashion. Sections are organized based on the main categories of visualization metrics in Sec. 4.3. The contents of a section in the report include both natural language and illustrative images (such as heatmaps, extracted shapes, or gaze maps) (see Fig. 1 (C)). Each section provides a short overall summary of suggestions (Fig. 1 (C) ②) of the underlying data for that metric followed by two standardized paragraphs:
  • Explanations: Explanation of the visualization feedback in natural language (“clarify” in the ACGT workflow) (e.g., Fig. 1 (C) ④).
  • Suggestions: Recommendations addressing shortcomings in the visualization for that specific metric (“guide”) (e.g., Fig. 1 (C) ③).
We provide additional functionality to help navigating the report. We indicate sections that require modification using a circular visual mark, either yellow or red next to the title of each section (see Fig. 1 (C) ①). If there exist at most one issue in a section, then the corresponding circle is yellow. If there is more than one, then the circle is orange. We provide a short overview summary of responses per category in Fig. 1 (E). This aids designers in quickly identifying their topics of interest by skimming section summaries before delving into specifics.
4.5 Revision Tracker
@@
The revision tracker gives a high-level overview of changes from one revision to another for the entire design process. This component realizes the “track” in the ACGT workflow. The intention is to give the designer a bird’s eye view of the design process, thereby seeing the improvements over time and also avoiding local optima. The natural language in the tracker is derived using the LLM with the reports to be compared as prompts; see below. For detailed tracking, Visualizationary allows designers to check all versions of their past reports by going to ‘Archives’ in the navigation bar (see Fig. 1 ⑧). The interface provides two separate screens, each displaying a report of a version, to facilitate the comparison process.
@@ 4.6 Prompt Engineering for Visualization Feedback
ACG-template. We call the analyze-clarify-guide (ACG) prompt template $T_{acg}$. [Q] is the question we define to create interpretations and suggestions. For instance, we use “Analyze the visual salience on text. Provide interpretations in 2 sentences.” or “Provide suggestions about the result of the previous question in 2 sentences.” We define 9–12 questions for each visualization metric. We replace the metric name in different questions, e.g., from visualization salience to color blindness. We call the prompt conditions [cond]: “Please interpret exactly in the following way, as if you are an assistant to a visualization designer, explaining to novice visualization designers. If no visual salience is detected, then just interpret it as No salience is detected in the chart image. If (the rate of salience in the textual zone over the rate of the textual zone in chart image) from the measured result is less than 0.6, then interpret it as a lack of salience in textual elements.” At the end of each prompt, we append [filter-suggestions], a compact representation of visualization design knowledge in natural language. [filter-suggestions] is available in the supplementary material.
@@ 4.7 Implementation Notes
The Visualizationary tool is implemented as a client-server web application using HTML, CSS, JavaScript, and JQuery. We use the Python Flask web framework [^2] https://flask.palletsprojects.com/; the analysis components were implemented server-side in Python 3. We store data into MongoDB. [^3] https://www.mongodb.com/ During the user study, Visualizationary was hosted on Amazon Web Services (AWS). We use gpt-3.5-turbo as LLM.
@@ 5 User Study
We performed a user study to assess the impact of LLM-generated feedback in the visualization design pipeline using the Visualizationary tool. We opted for a qualitative study where designers of different skill were asked to use Visualizationary to support the design process of a new visualization over several days. We then let a small group of visualization experts (with more than 10 years of experience in the field) assess the design process without seeing the Visualizationary reports. We also interviewed our participants to collect their perception of the impact of Visualizationary on their design work. The study received approval from our university’s ethics review board (IRB).
Table I: User study demographics. A total of 13 visualization designers of different seniority participated in the user study. These participants were all above the age of 18, had at least a bachelor’s degree in visualization or related field, and had more than three years of experience in designing visualizations. We categorize those with less than or equal to 3 years of experience in visualization as novice (6 participants), those with 4 to 7 years as intermediate (4 participants), and those with more than 7 years as expert groups (3 participants), respectively.
ID Gender Age Education Job Title Years in Vis Group
P1 Male 35 Ph.D. Postdoc in Visualization 11 years Expert
P2 Male 26 Bachelor’s Degree. Ph.D. student in Computer Science 3 years Novice
P3 Male 34 Ph.D. Research scientist 5 years Intermediate
P4 Male 29 Master’s Degree Software engineer 3 years Novice
P5 Male 28 Master’s Degree Ph.D. student in Computer Science 6 years Intermediate
P6 Male 31 Master’s Degree Ph.D. student in Computer Science 6 years Intermediate
P7 Male 30 Bachelor’s Degree Ph.D. student in Computer Science 5 years Intermediate
P8 Male 33 Ph.D. Assistant professor in Visualization 10 years Expert
P9 Male 31 Master’s Degree Data scientist 9 years Expert
P10 Female 24 Bachelor’s Degree Master’s student in Data Science 3 years Novice
P11 Male 33 Bachelor’s Degree Software engineer 1 year Novice
P12 Female 33 Bachelor’s Degree Primary school teacher 1 year Novice
P13 Male 32 Ph.D. Postdoc in Computer Science 1 year Novice
@@ 5.1 Participants
We advertised our study through messages to the Data Visualization Society, to alumni of our institution, and to other academic institutions. Out of many applicants, we selected only those who are over the age of 18, are capable of reading and speaking English fluently, and have experience in visualization for more than three years. As a result, we were able to recruit 13 visualization designers of different seniority. Table I shows the demographics of the participants. These participants were paid a \$40 Amazon gift card upon concluding the study.
While Visualizationary was originally designed for novice and intermediate visualization designers, our participants also included expert ones. We included such participants because we were interested in seeing whether our tool could benefit even an expert.
Beyond the designer participants, we also recruited three expert participants with at least five years of experience in visualization: an associate professor (or equivalent), an assistant professor, and a Ph.D. candidate (within months of graduation) within the research field. These experts only took part in the final independent design assessment step.
@@ 5.3 Procedure
Step 1: Pre-study Briefing. We started each session with an overview of our study, answered any questions from the participant, and collected informed consent using an online consent form, as per the directives of our IRB. Then the experiment administrator explained the longitudinal chart design task (Sec. 5.2). Next, the administrator showed how to operate the Visualizationary system: creating an account, logging in, uploading screenshots, and accessing design reports. The administrator answered any additional questions from the participant. The study was concluded by finding an appointment for the post-interview. On average, this briefing process took 30 minutes.
Step 2: Longitudinal Chart Design. Participants were given at least three days to design their visualization, during which time they were asked to upload at least five versions of their visualization. We imposed no restrictions on how participants allocated their time but asked that they dedicate at least a total of two hours during the period on the design and that they use the Visualizationary interface. During this period, the experiment administrator was available to answer technical questions.
@@
Step 4: Expert Design Assessment. In the final step, our independent expert participants (all with experience teaching data visualization and grading visualization assignments) were given all five designs for each of the designer participants. They were asked to provide their assessment of the quality of changes during the design process from the first to the final design for each participant using a 1–5 Likert scale (1 = significant decline in quality, 3 = neutral, 5 = significant quality improvement). They were also asked to motivate their assessment using 1 to 2 sentences. This meant that each expert gave 1 Likert scale rating and 1-2 assessment sentences per designer participant.

### v3

@@ 4.4 Feedback System
    Virtual Eyetracker. Virtual eyetracker refers to the extent to which areas within a scene capture the attention of an observer [55]. We utilize a virtual eyetracker [60] that predicts gaze on an overview task of a visualization image, and provide quantitative guidance. From the eyetracker, users can find where the audience would look at a visualization image as an overview. We have two reasons for providing metrics as guidance. First, we notify designers if heavy focus of salience at an area is indeed what they intended. As a case in point, if salience is primarily directed toward text, the metrics can prompt designers to reflect on whether they should shift more attention to graphical elements. Second, we alert users to potential biases by providing metrics that indicate when salience is heavily concentrated in a particular area. Because the eyetracker’s performance is not perfect, it is important to inform users about possible biases, particularly when salience rates are unusually high or low. For example, the center of an image may show unusually high salience even if the area contains no graphical elements.
    • A virtual eyetracker. Role: Predicts visual salience in a visualization to estimate what readers will notice during an overview task. Analysis process: A Scanner Deeply [60] generates an expected salience map about the inputted chart image on an overview task. Clarification: The salience map is overlaid as a heatmap on the chart image. The most salient areas appear in red or yellow, while less salient areas are shown in darker tones. Guidance to users: The salience map indicates whether the highlighted areas match the designer’s intended focus [10]. It also suggests methods to improve salience in specific zones. In order to help designers focus on their areas of interests, we introduce the concept of minimizing the data-ink ratio [65], methods to highlight zones of interest (e.g., using colors, making bold, etc.), and impact of grids in charts [69].
    • Focus on text. Role: Warns designers when the salience may focus overly on text. Analysis process: OCR software is used to detect textual zones, and texts are boxed by each letter. Then, we calculate how much salience is focused on these boxes. Clarification: Mark as overly salient to text if the salience in text boxes is above the $10^{th}$ percentile of the chart saliency in the dataset by Shin et al. [60]. Guidance to users: We warn designers that the salience of text is higher than other visualizations. We caution that the model’s salience measurements may be biased or inaccurate because it tends to highlight text too readily [60]. The designer has to decide whether to believe the results or not.
    • Focus on center. Role: Warns designers when the salience focuses overly on the center of the chart image. Analysis process: We measure the concentration of salience in the center. The center is defined as a rectangle in the middle of the image, with its height and width as 1/3×1/3$\times$1 / 3 × of the chart. Clarification: A visualization is flagged as overly salient in the center if saliency in this region is above $10^{th}$ percentile. Guidance to users: We warn designers that the salience on the central zone [9] is higher than other visualizations. We also inform the possibility that the model’s salience measurements may be biased or inaccurate.
    • Focus on visual attention. Role: Warns designers when a visualization may not effectively direct visual attention to important regions. Analysis process: Identify areas likely to attract attention by locating zones near RGB color transitions (within a 5-pixel radius), and measure how much of these zones are covered by high saliency values. Clarification: A visualization is flagged as “scarcely salient” if saliency in these regions falls below the 90th percentile of visualizations in the dataset. Guidance to users: Designers are alerted when attention-relevant zones receive less saliency than is typical. This matters because strong visualizations generally guide attention toward meaningful visual structures. Low saliency in such areas may reduce effectiveness.
@@
    Textual characters. Textual characters are visual objects that are used for labels, titles, and legends in visualization. Using text alongside charts can improve comprehension of visualizations [62]. To support this process, the system aims to detect title, and textual information in the chart.
    • Title. Role: Detects a title and provides recommendations. Analysis process: Title is detected via DePlot [35]. DePlot is a VLM that translates visualizations into tables and textual components. Gathering all the information, ChatGPT provides title suggestions for the chart. Clarification: DePlot reads the chart image, and decides if there is a title in the image. Guidance to users: If a title is not detected, then recommend adding a one and provide suggestions.
    • Textual content. Role: Detects text in visualization image. Analysis process: The OCR filter scans the chart image, and detects textual information. We use the PyTesseract-OCR as our OCR filter [64]. PyTesseract-OCR is a Python library that detects textual information from the visualization image. Clarification: If the OCR filter detects at least one letter in the text, then it contains the text. Guidance to users: When text is detected the system explains the benefits of texts in charts. If no text is detected, it explains in detail the benefits of having textual explanations in visualizations [62].
@@
    Visual representations. There exist various functionalities by knowing the chart’s data and its visual representations, such as visualization linting [13, 29]. Our focus is on providing chart recommendations based on the data table and design preamble, as well as detecting unnecessary visual embellishments (chartjunk). By leveraging these suggestions, we aim to improve the visual representation of the data.
    • Optimal chart type: Role: Suggests the designer an optimal chart, based on the data table translated from the chart. Analysis process: The data table is extracted from DePlot. The suggestion is generated based on a preamble of design knowledge and ChatGPT, which is described below. Clarification: The suggestions are determined based on two types of information: (1) the guideline as a preamble of design knowledge (perceptual rankings of visual cues, described in Mackinlay’s APT [40]) and (2) the data table provided by DePlot. Using these two information, ChatGPT makes the final suggestion. We recognize that this is a relatively simple model with limitations, such as the inability to recommend treemaps and a restricted knowledge base for certain visualization choices. No suggestion is provided if data table, or data is not extractable from the chart image. Guidance to users: A chart is suggested, and ChatGPT provides both the data table and the reasoning behind its choice. This helps the designer decide whether to follow the advice.
    • Visual embellishment (“chart junk”): Role: Detects any chartjunk [6] in the visualization, and if present, warns about its potential benefits and drawbacks. Analysis process: A computer vision model, such as YoloR [66] detects objects. Clarification: If the object detection model detects objects within the chart image, then we say that the model has chartjunk. Guidance to users: If chartjunk is detected, then warn of its benefits as well as dangers [7, 6].
@@
    Color perception. Color perception is a basic building block of data visualization [48, 54]. Visualization practice stipulates using a limited number of distinguishable and easily named colors [28], potentially as a function of the mark [63]. We detect the number of different and similar colors in a chart image. We then remind users about appropriate color usage. For instance, similar hues should represent continuous values, while distinct colors should signify categories [25, 48].
    • Variability of colors: Role: Detects different colors and reminds the designer to check if different colors are used effectively (representing categorical values). Analysis process: A custom function processes the chart image to identify distinct colors. Colors are grouped as similar if their maximum Euclidean distance in RGB space is less than 10, with each group represented by its centroid. Each group is treated as a single distinct color. Clarification: If the function identifies more than 2 distinct colors from the chart image, then the image is identified as having multiple colors. Guidance to users: After showing all distinct colors used, the filter recommends that different colors represent categories. However, it is not capable of detecting mismatch of colors.
    • Similarity of colors: Role: Detects similar colors and suggests the designer to identify if they are used effectively (representing continuous values). Analysis process: A customized function created in Python detects groups of similar colors (the definition is identical to similar colors described in variability of colors). A group of similar colors are determined using differences in RGB values between detected colors. Clarification: If the function identifies more than 2 similar groups of colors then the image is identified as having similar colors. Guidance to users: The filter recommends that similar colors represent continuous, related values. This filter too, is not capable of detecting color mismatch.
@@
    Accessibility. Accessibility is emerging as an important new topic in the visualization field [21], including issues such as color vision deficiency, blindness, and low vision. [19, 15, 17, 42, 21]. Although accessibility encompasses many considerations, we specifically addresses color vision deficiency (CVD). We focus on CVD because it affects a significant portion of the population (about 1 in 12 males and 1 in 200 females are affected by CVD) and directly impacts the effectiveness of color-encoded data [57].
    • CVD: Role: Helps detect visualizations that are significantly affected for people with color vision deficiency (deuteranopia, protanopia, tritanopia). Analysis process: We evaluate the chart image from two perspectives. First, we quantitatively measure the information loss in the simulated visualizations using the entropy difference from the original image. Second, we simulate the chart image as if it were viewed with color vision deficiency. Clarification: If the entropy loss in the simulated visualization image exceeds a certain threshold, then the chart image is marked as significantly affected. Guidance to users: We caution that individuals with color vision deficiency may miss information. To observe the exact differences, one should view the simulated chart image. We then provide recommended CVD-safe palettes. However, it is up to the designer to select the palette.
When the system detects no quantitative issue in clarification, it indicates that no problem was found and then presents the filter’s intent. For example, when a title is detected, the system says, “A title is detected. The intention of this filter is to determine whether a title exists and, if it does not, recommend adding one, because a title helps users better understand the visualization.” Another example is when the system detects no chartjunk. In that case, the system replies as, “We did not detect any chartjunk.
@@
Revision Tracker. The revision tracker in Fig. 2 (D) gives a high-level overview of changes from one revision to another for the entire design process. This component realizes the “track” in the ACGT workflow. The intention is to give the designer a bird’s eye view of the design process, thereby seeing the improvements over time and also avoiding local optima. The natural language in the tracker is derived using the LLM with the reports to be compared as prompts; see the next section. The revision tracker primarily monitors changes in quantifiable metrics for each topic. It displays any increases or decreases and provides commentary on whether those changes are beneficial. Because it focuses on metrics, it does not capture qualitative comparisons (e.g., saliency maps or simulated CVD images). To address that gap, we also offer a detailed tracking capability. For detailed tracking, Visualizationary allows designers to check all versions of their past reports by going to ‘Archives’ in the navigation bar (see Fig. 2 ⑧). The interface provides two separate screens, each displaying a report of past version selected by the user, to facilitate the comparison process.
@@ 4.5 Prompt Engineering for Visualization Feedback
To further minimize hallucinations, we incorporate relevant grounding contexts (e.g., design knowledge about data visualizations and filters) [44] in the prompt, preventing the LLM from generating text outside its learned scope.
@@
ACG-template. We call the analyze-clarify-guide (ACG) prompt template $T_{acg}$. [Q] is the question we define to create interpretations and suggestions. For instance, we use “Analyze the visual salience on text. Provide interpretations in 2 sentences.” or “Provide suggestions about the result of the previous question in 2 sentences.” We define 9–12 questions for each visualization metric. We replace the metric name in different questions, e.g., from visualization salience to color blindness. We call the prompt conditions [cond]: “Please interpret exactly in the following way, as if you are an assistant to a visualization designer, explaining to novice visualization designers. If no visual salience is detected, then just interpret it as No salience is detected in the chart image. If (the rate of salience in the textual zone over the rate of the textual zone in chart image) from the measured result is less than 0.6, then interpret it as a lack of salience in textual elements.” At the end of each prompt, we append [filter-suggestions], a compact representation of visualization design knowledge in natural language; see the supplementary material.
We mainly use “if, then” tone when providing design knowledge. While this may not be the most natural style of discussion, we want to convey that our suggestions are optional and to prevent designers from applying feedback inaccurately.
@@ 4.6 Implementation Notes
The Visualizationary system is implemented as a web application using HTML, CSS, JavaScript, and JQuery. We use the Python Flask web framework [^2] https://flask.palletsprojects.com/; the analysis components were implemented server-side in Python 3. We store data into MongoDB. [^3] https://www.mongodb.com/ During the user study, Visualizationary was hosted on Amazon Web Services (AWS). We use gpt-3.5-turbo as LLM.
@@ 5 User Study
We performed a user study to assess the impact of LLM-generated feedback in the visualization design pipeline using the Visualizationary system. We opted for a qualitative study where designers of different skill were asked to use Visualizationary to support the design process of a new visualization over several days. We then let a small group of visualization experts assess the design process without seeing the Visualizationary reports. We also interviewed our participants to collect their perception of the impact of Visualizationary on their design work. The study received approval from our university’s ethics review board (IRB).
Table I: User study demographics. A total of 13 visualization designers of different seniority participated in the user study. These participants were all above the age of 18, had at least a bachelor’s degree in visualization or related field, and had more than a year of experience in designing visualizations. We categorize those with less than or equal to 3 years of experience in visualization as novice (6 participants, N1-N6), those with 4 to 7 years as intermediate (4 participants, I1-I4), and those with more than 7 years as expert groups (3 participants, E1-E3), respectively. We present participants’ gender, age, education (Edu), job title, and years of experience in Vis (Exp.).
ID Gender Age Edu. Job Title Exp.
N1 Male 33 B.S. Software engineer 1 yr
N2 Female 33 B.S. Primary school teacher 1 yr
N3 Male 32 Ph.D. Postdoc in CS 1 yr
N4 Male 26 B.S. Ph.D. student in CS 3 yrs
N5 Male 29 M.S. Software engineer 3 yrs
N6 Female 24 B.S. M.S. student in Data Science 3 yrs
I1 Male 30 B.S. Ph.D. student in CS 5 yrs
I2 Male 34 Ph.D. Research scientist 5 yrs
I3 Male 28 M.S. Ph.D. student in CS 6 yrs
I4 Male 31 M.S. Ph.D. student in CS 6 yrs
E1 Male 31 M.S. Data scientist 9 yrs
E2 Male 33 Ph.D. Assistant professor in Vis 10 yrs
E3 Male 35 Ph.D. Postdoc in Vis 11 yrs
Table II: Participants’ study setup. Here we list the settings the participants used to conduct our study. We present the tools, datasets and chart types used by the participants. We also show the total time (hh:mm) taken for these participants to finish the work (all participants updated five designs within a single continuous time frame).
ID Tool Used Dataset Topics Chart Type Explored Time Between V1 and V5
N1 Microsoft Excel Time management Stacked area/bar charts, line chart 01:47
N2 Microsoft Excel Survey analysis Bar chart 01:23
N3 Python (Matplotlib) Topic document analysis Bar chart 01:14
N4 Python (Matplotlib) Scientific result analysis Bar chart, line chart 02:16
N5 Python (Matplotlib) Dataset analysis Line chart, bar chart 01:48
N6 Tableau Scientific result analysis Dot plot, line chart 02:31
I1 Javascript (D3.js) Cluster analysis Matrix scatterplot 01:54
I2 Tableau Credit card usage Bar chart, Stacked bar chart 02:28
I3 Microsoft Excel User experience on UI Bar chart 03:18
I4 Microsoft Excel Scientific result analysis Bar chart, stacked bar chart 02:04
E1 Tableau Sales data analysis Geographic map, bubble chart 02:35
E2 Spotfire Market analysis Bar chart 02:15
E3 Microsoft Excel Scientific result analysis Boxplot 01:47
@@ 5.1 Participants
We advertised our study through messages to the Data Visualization Society, to alumni of our institution, and to other academic institutions. Out of many applicants, we selected only those who are over the age of 18, are capable of reading and speaking English fluently, and have experience in visualization for more than a year. As a result, we were able to recruit 13 visualization designers of different seniority. Table I shows the demographics of the participants. These participants were paid a \$40 Amazon gift card upon concluding the study. While Visualizationary was originally designed for novice and intermediate visualization designers, our participants also included expert ones. We included such participants because we were interested in seeing whether our system could benefit even an expert.
Beyond the designer participants, we also recruited three expert participants with at least seven years of experience in visualization: an associate professor (or equivalent), an assistant professor, and a Ph.D. candidate (within months of graduation) within the research field. These experts only took part in the final independent design assessment step.
@@ 5.3 Procedure
Step 1: Pre-study Briefing. We started each session with an overview of our study, answered any questions from the participant, and collected informed consent using an online consent form, as per the directives of our IRB. Then the experiment administrator explained the longitudinal chart design task (§ 5.2). Next, the administrator showed how to operate the Visualizationary system: creating an account, logging in, uploading screenshots, and accessing design reports. The administrator answered any additional questions from the participant. The study was concluded by finding an appointment for the post-interview. On average, this briefing process took 30 minutes.
Step 2: Longitudinal Chart Design. Participants were given at least three days to design their visualization, during which time they were asked to upload at least five versions of their visualization. We imposed no restrictions on how participants allocated their time (R2) but asked that they dedicate at least a total of two hours during the period on the design and that they use the Visualizationary interface (R1). During this period, the experiment administrator was available to answer technical questions.
@@
Step 4: Expert Design Assessment. In the final step, our independent expert participants (all with experience teaching data visualization and grading visualization assignments) were given all five designs for each of the designer participants. They were asked to provide their assessment of the quality of changes during the design process from the first to the final design for each participant using a 1–5 Likert scale (1 = significant decline in quality, 3 = neutral, 5 = significant quality improvement). They were also asked to motivate their assessment using 1 to 2 sentences. This meant that each expert gave 1 Likert scale rating and 1-2 assessment sentences per designer participant’s each iteration (that is, 5 ratings and 5 assessment sentences in total).